OpenStudy (anonymous):

SHOW THAT THERE ARE NO TWO INTEGERS X AND Y THAT SOLVE X^2 +3Y^2=8

OpenStudy (anonymous):

Integers are whole numbers therefore looking at the equation solved for either variable neither y=sqrt((8-x^2/3) or x=sqrt(8-3y^2) will never equal a whole numbers if x and y are limited to integer values so they will never =8... probably an easier way to say that but i think that works